# Magnon squeezing near a quantum critical point in a cavity-magnon-qubit system

### v1

@@ Abstract
Preparing magnon nonclassical states is a central topic in the study of quantum magnonics. Here we propose to generate magnon squeezed states in a hybrid cavity-magnon-qubit system by engineering an effective Rabi-type magnon-qubit interaction. This is achieved by adiabatically eliminating the cavity mode and driving the qubit with two microwave fields, of which the driving frequencies and amplitudes are properly selected. By operating the system around the critical point associated with the ground-state superradiant phase transition in the normal phase, a magnon parametric amplification-like interaction is induced, leading to a dynamical magnon squeezing. We further analyze the effects of the dissipation, dephasing, and thermal noise on the magnon squeezing. Our results indicate that a moderate degree of squeezing can be produced using currently available parameters in the experiments.
@@ I INTRODUCTION
In recent years, a number of proposals have been offered to prepare magnon squeezed states, e.g., by transferring squeezing from a microwave field Li et al. (2019) or a phonon mode Fan et al. (2024) to the magnon mode, two-tone driving of the magnon mode Zhang et al. (2021); Qian et al. (2024), exploiting the anisotropy of the ferromagnet Kamra and Belzig (2016) and the mechanism of the ponderomotivelike squeezing Li et al. (2022), and using an easy-axis ferromagnet Lee et al. (2023). In cavity-magnon-qubit systems, magnon squeezing can be achieved by two-tone driving of a superconducting qubit to induce the magnonic parametric amplification Guo et al. (2023), and engineering a nonlinear magnon-qubit coupling Xia et al. (2025).
Here, we propose a different mechanism for achieving magnon squeezing by engineering a Rabi-type magnon-qubit interaction in the cavity-magnon-qubit system. Specifically, when the microwave cavity is far detuned from both the magnon and qubit system, the cavity mode can be adiabatically eliminated, yielding an effective magnon-qubit coupling mediated by virtual photons. By driving the superconducting qubit with two microwave fields and appropriately adjusting the drive frequencies and amplitudes, we show that the magnon-qubit system can realize an effective quantum Rabi model Ballester et al. (2012); Braumüller et al. (2017). Notably, when operating the system approaching the critical point associated with the ground-state superradiant phase transition in the normal phase, a parametric amplification-like interaction arises for the magnon mode, which facilitates the dynamical generation of a magnon squeezed state via a two-magnon process. Using experimentally accessible parameters, our scheme can achieve a magnon squeezing of about 4 dB. The generated squeezing is robust against both the system dissipations and thermal noise. Compared to the existing qubit-assisted schemes Guo et al. (2023); Xia et al. (2025), our approach features a distinct underlying mechanism. In Ref. Guo et al. (2023), the cavity mediates an effective Jaynes-Cummings (JC)-type interaction between the qubit and the magnon mode, and two-tone driving of the qubit induces the magnonic parametric amplification. In Ref. Xia et al. (2025), virtual cavity photons generate an effective nonlinear coupling that leads to a two-magnon interaction and hence magnon squeezing. By contrast, our scheme engineers an effective Rabi-type magnon-qubit interaction, where operating near the critical point enables criticality-enhanced parametric amplification, leading to a moderate and tunable magnon squeezing.
@@ II The system and the effective magnon-qubit Hamiltonian
where $\omega_{m}=\omega_{M}+g_{\rm cm}^{2}/\Delta_{m}$ is the effective frequency of the magnon mode, $\omega_{q}=\omega_{Q}+g_{\rm cq}^{2}/\Delta_{q}$ is the effective frequency of the superconducting qubit, as depicted in Fig. 1(b), and $G=\frac{g_{\rm cq}g_{\rm cm}}{2}(\Delta_{m}^{-1}+\Delta_{q}^{-1})$ is the effective coupling strength between the qubit and the magnon mode.
@@
with $g=G/2$. This Hamiltonian can capture the dynamics and rich physics in the ultrastrong ($0.1<g/\delta_{m}<1$) and even deep-strong ($g/\delta_{m}\gtrsim 1$) coupling regimes Frisk Kockum et al. (2019); Forn-Díaz et al. (2019).
@@
Figure 2: (a) The minimum variance of the quadrature $V_{\rm min}(X_{\theta})$ and (b) the degree of squeezing $S$ (dB) as a function of time, which are obtained at an optimal squeezing angle at each time point. In (a) and (b), the black curves correspond to the results obtained with the full Hamiltonian in Eq. (II), while the red thinner curves are obtained using the approximate Hamiltonian in Eq. (7). (c) The optimal squeezing angle $\theta_{\rm opt}$ and (d) the mean magnon number $\braket{m^{\dagger}m}$ during the evolution, which are obtained using the master equation in Eq. (III). The parameters adopted are provided in the main text.
@@ III Dynamical magnon squeezing
In Figs. 2(a) and 2(b), we plot the minimum variance $V_{\rm min}(X_{\theta})$ and the degree of squeezing $S$ as a function of time, and compare the results obtained using the derived Rabi-type Hamiltonian in Eq. (7) with those using the full Hamiltonian in Eq. (II). The results indicate that a magnon squeezing of approximately 4 dB can be achieved at an optimal time, and the Hamiltonian (7) is a good approximation especially at the time corresponding to the maximum squeezing. We use the following experimentally feasible parameters Tabuchi et al. (2015); Lachance-Quirion et al. (2017); Xu et al. (2023): $\omega_{c}/2\pi=6.2180\,\mathrm{GHz}$, $\omega_{Q}/2\pi=5.8447\,\mathrm{GHz}$, $\omega_{M}/2\pi=5.9205\,\mathrm{GHz}$, $g_{\rm cq}/2\pi=74.7\,\mathrm{MHz}$ and $g_{\rm cm}/2\pi=59.5\,\mathrm{MHz}$, which lead to the detunings $\Delta_{q}/2\pi=373.3\,\mathrm{MHz}$ and $\Delta_{m}/2\pi=297.5\,\mathrm{MHz}$, the effective frequencies $\omega_{q}/2\pi=5.8596\,\mathrm{GHz}$ and $\omega_{m}/2\pi=5.9324\,\mathrm{GHz}$, and the effective coupling strength $G/2\pi=13.4\,\mathrm{MHz}$. For the drives, we set $\mathcal{E}_{1}/2\pi=500~\text{MHz}$, $\mathcal{E}_{2}/2\pi=60~\text{MHz}$, $\omega_{1}/2\pi=5.9294~\text{GHz}$, $\omega_{2}/2\pi=4.9294~\text{GHz}$, and $\delta_{m}/2\pi=3.0~\text{MHz}$. The other parameters are $\kappa/2\pi=0.5$ MHz Shen et al. (2025), $\gamma/2\pi=\gamma_{\phi}/2\pi=3$ kHz, and $T=10$ mK. We note that by using a stronger drive field, yielding a larger $\mathcal{E}_{1}$, the condition for taking the RWA to get Eq. (7) is better satisfied, which will lead the two curves in Figs. 2(a) and 2(b) to be closer.
@@
To give a full picture of the squeezed state generated in our scheme, we plot in Fig. 4(b) the Wigner function of the magnon mode corresponding to the maximum squeezing at an optimal time, defined as
$W(\alpha,\alpha^{*})=\frac{2}{\pi}\text{Tr}\left[e^{i\pi m^{\dagger}m}D(\alpha){\rho}D^{\dagger}(\alpha)\right],$ (14)
which serves as a quasi-probability distribution providing a complete phase-space representation of the state. Here, $D(\alpha)=\exp(\alpha m^{\dagger}-\alpha^{*}m)$ is the displacement operator for the magnon mode with $\alpha$ being a complex variable. Figure 4(b) clearly exhibits a squeezed magnon mode at an optimal squeezing angle.
@@ IV Conclusions
In summary, we have presented an experimentally feasible scheme for generating magnon squeezed states in a microwave cavity-magnon-qubit system. By precisely controlling the frequencies and amplitudes of the two driving fields applied to the qubit, an effective Rabi-type magnon-qubit interaction can be achieved, which leads to a dynamical magnon squeezed state when operating the system around the critical point associated with the ground-state phase transition. Our results indicate that a moderate degree of magnon squeezing about $4$ dB can be achieved using currently available parameters in relevant experiments, which highlight the potential of the qubit-assisted hybrid platforms for engineering magnon nonclassical states. The magnon squeezed states offer promising opportunities for quantum information processing and quantum sensing with magnons.

### v2

@@ Abstract
Preparing magnon nonclassical states is a central topic in the study of quantum magnonics. Here we propose to generate magnon squeezed states in a hybrid cavity-magnon-qubit system by engineering an effective Rabi-type magnon-qubit interaction. This is achieved by adiabatically eliminating the cavity mode and driving the qubit with two microwave fields, of which the driving frequencies and amplitudes are properly selected. In the normal phase of the effective Rabi model associated with the ground-state superradiant phase transition, the driven magnon–qubit dynamics generate a parametric-amplification-like two-magnon interaction that produces magnon squeezing. The resulting squeezing is strongly enhanced when the system is tuned close to the critical point. We further analyze the effects of the dissipation, dephasing, and thermal noise on the magnon squeezing. Our results indicate that a moderate degree of squeezing can be produced using currently available parameters in the experiments.
@@ I INTRODUCTION
In recent years, a number of proposals have been offered to prepare magnon squeezed states, e.g., by transferring squeezing from a microwave field Li et al. (2019) or a phonon mode Fan et al. (2024) to the magnon mode, two-tone driving of the magnon mode Zhang et al. (2021); Qian et al. (2024), exploiting the anisotropy of the ferromagnet Kamra and Belzig (2016) and the mechanism of the ponderomotivelike squeezing Li et al. (2022), and using an easy-axis ferromagnet Lee et al. (2023). In cavity-magnon-qubit systems, magnon squeezing can be achieved by two-tone driving of a superconducting qubit to induce the magnonic parametric amplification Guo et al. (2023), and engineering a nonlinear magnon-qubit coupling Xia et al. (2025). We note that the pulse protocol Sharma et al. (2022) has a potential to prepare a magnon squeezed state, but due to the very short magnon lifetime and limited magnon-qubit coupling strength in the practical situation Lachance-Quirion et al. (2020); Xu et al. (2023), the protocol is limited to generate a superposition state of Fock states $|N\rangle$ up to $N=3$, which is incompetent to construct a squeezed vacuum state that corresponds to an infinite dimensional Hilbert space.
Here, we propose a different mechanism for achieving magnon squeezing by engineering a Rabi-type magnon-qubit interaction in the cavity-magnon-qubit system. Specifically, when the microwave cavity is far detuned from both the magnon and qubit system, the cavity mode can be adiabatically eliminated, yielding an effective magnon-qubit coupling mediated by virtual photons. By driving the superconducting qubit with two microwave fields and appropriately adjusting the drive frequencies and amplitudes, we show that the magnon-qubit system can realize an effective quantum Rabi model Ballester et al. (2012); Braumüller et al. (2017). Notably, in the normal phase the effective magnon–qubit dynamics already contain a parametric-amplification-like two-magnon interaction for the magnon mode. This interaction generates magnon squeezing and becomes increasingly strong as the system is tuned towards the critical point associated with the ground-state superradiant phase transition. Using experimentally accessible parameters, our scheme can achieve a magnon squeezing of about 3.7 dB. The generated squeezing is robust against both the system dissipations and thermal noise. Compared to the existing qubit-assisted schemes Guo et al. (2023); Xia et al. (2025), our approach features a distinct underlying mechanism. Specifically, Ref. Guo et al. (2023) achieves magnon squeezing through direct parametric amplification within a driven effective Jaynes-Cummings (JC) model. By contrast, our scheme critically relies on the phase transition of an effective quantum Rabi model. The squeezing mechanism of the Ref. Xia et al. (2025) is based on an engineered nonlinear coupling between the superconducting qubit and the microwave resonator, which is however only present for a microwave transmission line resonator but absent for a 3D microwave cavity.
@@ II The system and the effective magnon-qubit Hamiltonian
where $\omega_{m}=\omega_{M}+g_{\rm cm}^{2}/\Delta_{m}$ is the effective frequency of the magnon mode, $\omega_{q}=\omega_{Q}+g_{\rm cq}^{2}/\Delta_{q}$ is the effective frequency of the superconducting qubit, as depicted in Fig. 1(b), and $G={g_{\rm cq}g_{\rm cm}}(\Delta_{m}^{-1}+\Delta_{q}^{-1})/2$ is the effective coupling strength between the qubit and the magnon mode.
@@
with $g=G/2$. The validity of the RWA used to derive Eq. (7) is explicitly analysed in Appendix B. This Hamiltonian can capture the dynamics and rich physics in the ultrastrong ($0.1<g/\delta_{m}<1$) and even deep-strong ($g/\delta_{m}\gtrsim 1$) coupling regimes Frisk Kockum et al. (2019); Forn-Díaz et al. (2019).
@@
Figure 2: (a) The minimum variance of the quadrature $V_{\rm min}(X_{\theta})$ and (b) the degree of squeezing $S$ (dB) as a function of time, which are obtained at an optimal squeezing angle at each time point. In panels (a) and (b), the black curves are the results obtained from the full Hamiltonian in Eq. (II) including the two-tone driving, the blue thinner curves are obtained from Eq. (II) after adiabatic elimination of the cavity in the large-detuning regime, and the red thinnest curves correspond to the Hamiltonian in Eq. (7) after the RWA. (c) The optimal squeezing angle $\theta_{\rm opt}$ and (d) the mean magnon number $\braket{m^{\dagger}m}$ during the evolution, which are obtained using the master equation in Eq. (III). The parameters adopted are provided in the main text.
@@ III Dynamical magnon squeezing
In Figs. 2(a) and 2(b), we plot the minimum variance $V_{\rm min}(X_{\theta})$ and the degree of squeezing $S$ as a function of time, and compare the results obtained from the full Hamiltonian in Eq. (II) including the two-tone driving, the effective JC Hamiltonian in Eq. (II) after adiabatic elimination of the cavity, and the effective Rabi-type Hamiltonian in Eq. (7) after the RWA. The three models predict qualitatively consistent squeezing dynamics with almost identical time to achieve the maximum squeezing. Their moderate quantitative difference is due to the fundamental inconsistency between the requirements on system parameters for satisfying the condition of the approximation and for achieving a large degree of squeezing. Specifically, enforcing a strict large-detuning condition improves the accuracy of the adiabatic elimination and the RWA, but yields a reduced effective coupling strength $G$ and thus a weaker squeezing. Relaxing this condition enhances the nonlinear interaction and thus the degree of squeezing but at the cost of the deviation of the dynamics predicted by the three models. The results indicate that a magnon squeezing of approximately 3.7 dB can be achieved at an optimal time, and the Hamiltonian (7) is a good approximation especially at the time corresponding to the maximum squeezing. We use the following experimentally feasible parameters Tabuchi et al. (2015); Lachance-Quirion et al. (2017); Xu et al. (2023): $\omega_{c}/2\pi={6.4720}\,\mathrm{GHz}$, $\omega_{Q}/2\pi={5.8447}\,\mathrm{GHz}$, $\omega_{M}/2\pi={5.9205}\,\mathrm{GHz}$, $g_{\rm cq}/2\pi={85.7}\,\mathrm{MHz}$ and $g_{\rm cm}/2\pi={72.5}\,\mathrm{MHz}$, which lead to the detunings $\Delta_{q}/2\pi={627.3}\,\mathrm{MHz}$ and $\Delta_{m}/2\pi={551.5}\,\mathrm{MHz}$, the effective frequencies $\omega_{q}/2\pi={5.85641}\,\mathrm{GHz}$ and $\omega_{m}/2\pi={5.93003}\,\mathrm{GHz}$, and the effective coupling strength $G/2\pi={10.5854}\,\mathrm{MHz}$. For the drives, we set $\mathcal{E}_{1}/2\pi=500~\text{MHz}$, $\mathcal{E}_{2}/2\pi={50}~\text{MHz}$, $\omega_{1}/2\pi={5.928}~\text{GHz}$, $\omega_{2}/2\pi={4.928}~\text{GHz}$, and $\delta_{m}/2\pi={2.03083}~\text{MHz}$. The other parameters are $\kappa/2\pi=0.5$ MHz Shen et al. (2025), $\gamma/2\pi=\gamma_{\phi}/2\pi=3$ kHz, and $T=10$ mK. We note that by using a stronger drive field, yielding a larger $\mathcal{E}_{1}$, the condition for taking the RWA to get Eq. (7) is better satisfied, which will lead the two curves in Figs. 2(a) and 2(b) to be closer.
@@
In the experiment, the squeezing can be verified by performing the Wigner tomography of the magnon mode Xu et al. (2023). It is essentially to measure the expectation value of the magnon displaced parity operator, i.e.,
$W(\alpha)=\frac{2}{\pi}\text{Tr}\left[e^{i\pi m^{\dagger}m}D^{\dagger}(\alpha){\rho}D(\alpha)\right],$ (14)
where $D(\alpha)=\exp(\alpha m^{\dagger}-\alpha^{*}m)$ denotes the magnon displacement operator with $\alpha$ being a complex variable and $e^{i\pi m^{\dagger}m}=(-1)^{m^{\dagger}m}$ is the magnon parity operator. In Fig. 4(b), we plot the Wigner function of the magnon mode corresponding to the maximum squeezing at an optimal time, which clearly shows a squeezed magnon mode at an optimal squeezing angle.
@@ IV Conclusions
In summary, we have presented an experimentally feasible scheme for generating magnon squeezed states in a microwave cavity-magnon-qubit system. By precisely controlling the frequencies and amplitudes of the two driving fields applied to the qubit, an effective Rabi-type magnon-qubit interaction can be achieved, which yields a parametric-amplification-like two-magnon process that dynamically generates magnon squeezed states in the normal phase. The resulting squeezing can be strongly enhanced when the system is operated close to the ground-state phase-transition point. Our results indicate that a moderate degree of magnon squeezing about 3.7{3.7} dB can be achieved using currently available parameters in relevant experiments, which highlight the potential of the qubit-assisted hybrid platforms for engineering magnon nonclassical states. The magnon squeezed states offer promising opportunities for quantum information processing and quantum sensing with magnons.
@@ Acknowledgements.
Figure 6: Maximum degree of squeezing versus Rabi frequency $\mathcal{E}_{1}/2\pi$. The red dashed curve is obtained by using the full driven JC Hamiltonian Eq. (II), while the black solid line corresponds to the effective $\mathcal{E}_{1}$-independent Rabi model Eq. (7).
@@ Appendix A The effect of the relative phase between two driving fields
Appendix B Analysis of the validity of the RWA for the effective Rabi model
Here, we numerically assess the validity of the RWA used to derive the effective Rabi Hamiltonian Eq. (7) from the full driven JC model Eq. (II). Specifically, in Fig. 6 we numerically solve the master equation Eq. (III) for the JC Hamiltonian Eq. (II) at different values of the Rabi frequency $\mathcal{E}_{1}$ and achieve the maximum degree of squeezing at an optimal time. We then compare the results with those obtained by solving the master equation Eq. (III) for the effective Rabi Hamiltonian Eq. (7). Clearly, the approximation, i.e., the Rabi model, is more accurate as $\mathcal{E}_{1}$ increases, which better satisfies the condition of the RWA. Even for a relatively small value $\mathcal{E}_{1}/2\pi=200$ MHz, the deviation of the degree of squeezing is less than 1 dB, i.e., 4.353.55=0.84.35-3.55=0.8 dB. In plotting Fig. 6, we only vary the Rabi frequency $\mathcal{E}_{1}$ and accordingly the drive frequency $\omega_{2}$, such that $\omega_{1}-\omega_{2}=2\mathcal{E}_{1}$ is always kept. The other parameters are the same as those in Fig. 2.